Convert this equation -6x+3y=10 to slope intercept form and then identify the correct slope of this line.
Answer by rmromero(383)

-6x+3y=10 to slope intercept form and then identify the correct slope of this line.
 
The equation y = mx+b is called the Slope -intercept form  of a linear equation. 
The graph of the equation is a line with slope m and y-intercept b.

-6x + 3y = 10 
      3y = 6x + 10   Divide each term by 3
       y = 2x + 5/3  where m = 2 and b = 5/3

The slope = m = 2